OpenStudy (anonymous):

RS=6, RT=3, and TS=4 1/2. Find MR.

OpenStudy (anonymous):



OpenStudy (e.mccormick):

Well, from the given can you determine anything else?

OpenStudy (anonymous):

Umm .like I'm not sure if from the given I need everything in it? I know it might be like two seperate equations

OpenStudy (anonymous):

Because TS:RS = TS:SM right?

OpenStudy (e.mccormick):

If you know they are similar triangles, then yes, you can use rations to sove it.

OpenStudy (anonymous):

Well like the whole section of the book is on similar triangles .

OpenStudy (anonymous):

This isn't a proof so I think it's assumed they're similar ?

OpenStudy (e.mccormick):

Well, you can probably show they are using the given angle that is congruent.

OpenStudy (e.mccormick):

But lets just say they are. Then yes, you can make a ratio between them and find all the sides.

OpenStudy (anonymous):

So first I need to solve that. Like\[\frac{ 6 }{ 4.5 } = \frac{ SM }{ 4.5 }\]

OpenStudy (anonymous):

Oops wait.

OpenStudy (anonymous):

\[\frac{ 6 }{ 4.5 } = \frac{ 4.5 }{ SM }\]

OpenStudy (anonymous):

because ST is mean proportional ?

OpenStudy (e.mccormick):

Yah, that should work. And once you know the whole side, subtract off the excess.

OpenStudy (anonymous):

That would be 6SM=20.25? then divide ?

OpenStudy (e.mccormick):

Hmmm.... but it gives an odd number.

OpenStudy (anonymous):

Maybe turn it into a fraction\[6SM = 20 \frac{ 1}{ 4 }\]

OpenStudy (anonymous):

\[6SM = \frac{ 81 }{ 4 } , SM = \frac{ 81 }{ 24 }\]

OpenStudy (anonymous):

Now we need to do the proportion of TR?

OpenStudy (e.mccormick):

Note: It is the longest side. Therefore it must be more than 6 because 6 is part of it....

OpenStudy (anonymous):

Okay but now we do this? \[\frac{ MR }{ 3 } = \frac{ 3 }{ 6 }\]

OpenStudy (e.mccormick):

No. You have found MS=3.375. That is impossible.

OpenStudy (anonymous):

lol. idk. Maybe you can assume the picture to be like how the numbers are?

OpenStudy (e.mccormick):

Well, I am not sure you have identified what is similar properly. That makes it hard to get the right answer.

OpenStudy (anonymous):

Is not TM the bisector of angle PTR? Idk if that helps

OpenStudy (e.mccormick):

Yes, it is... hmmm....

OpenStudy (anonymous):

I'm looking up notes seeing if I can figure this with you

OpenStudy (anonymous):

a theorem says the bisector of exterior angle of triangles divides opposite side externally into segment which have some ratio as other 2 sides.

OpenStudy (e.mccormick):

Ah, found a diragram. You may have a direct ratio: http://jwilson.coe.uga.edu/emt725/Bisect/bisect.html

OpenStudy (e.mccormick):

On that one, it shows constructing: \(\dfrac{AD}{BD}=\dfrac{AC}{BC}\) Now, if you translate that to yours...

OpenStudy (e.mccormick):

OK, so used that info, or another version here: http://jwilson.coe.uga.edu/MATH7200/SupplementalTheorems/SupplementalTheorems.html to come up with this: \(\dfrac{TS}{RT}=\dfrac{MS}{MR}\) Now, what is known? \(\dfrac{4.5}{3}=\dfrac{MS}{MR}\) Does not look too good, two unknowns... BUT! MS = MR + RS MS = MR + 6 \(\dfrac{4.5}{3}=\dfrac{MR + 6}{MR}\) One unknown!

OpenStudy (e.mccormick):

Check the way I put it together with your "bisector of exterior angle of triangles divides opposite side externally into segment which have some ratio as other 2 sides" notes. You should be able to do that or something very similar. I may have a number in the wrong place... so check it! But that will lead to an answer.

OpenStudy (e.mccormick):

So I am guessing you got it.